If a+b+c=0,then the family of lines 3ax + by + 2c = 0 pass through the fixed point

Grade:12

1 Answers

Vikas TU
7 years ago
Dividing the whole eqn.   by 2 and comparing with a+b+c = 03ax + by + 2c = 0 
we get,
3a/2 x  +  b/2  y   +   c  = 0
 
3x/2 = 1
x = 2/3
and
y/2 = 1
y = 2.
hence (x,y) that is (2/3 , 2) is the fixed point .